How do you find a quadratic equation if you are only give the solution? Is it possible to have different quadratic equations with the same solution? Explain. give an example.
Thanks,
ts

Answer by stanbon(75887)

How do you find a quadratic equation if you are only give the solution? Is it possible to have different quadratic equations with the same solution? Explain. give an example.
------------------------
If "a" and "b" are solutions then the quadratic is
f(x) = k(x-a)(x-b)
Since "k" can take any Real number value except zero, there are many
equations that have the same solutions:
Ex:
y = 3(x-1)(x+5)
has the same solutions as
y = 7(x-1)(x+5)
=================
cheers,
Stan H.